# Quartet: Native FP4 Training Can Be Optimal for Large Language Models

### v3

@@ A.2 Scaling Law fitting
Identical to prior work [busbridge2025distillationscalinglaws], we fit the unquantized scaling law of the form
@@
We additionally for the scaling law forms with fixed $\gamma=1$ [hoffmann2022trainingcomputeoptimallargelanguage] and $\beta=1$ [kaplan2020scalinglawsneurallanguage]. The fits are presented in Figure 5 alongside the mainly used of busbridge2025distillationscalinglaws.
@@ A.3 Performance breakdown
Figure 6 presents a breakdown of runtime composition across three linear layer shapes in a Llama-7B model, taking the MXFP4 forward pass as an example. Each subplot shows the percentage of total runtime spent in three key kernel stages: matrix multiplication, quantization-related operations, and rearrangement of scaling factors for the mma instruction [Blackwell].
@@
As the figure illustrates, tuning the quantization kernel significantly reduces the proportion of time spent in the quantization stage—particularly for large matrix shapes. Increasing the threadblock tile size leads to more active warps per block, enhancing arithmetic intensity and enabling better latency hiding. In CUTLASS-based implementations, this change influences the multilevel tiling strategy (threadblock, warp, and instruction-level tiling), which is designed to optimize data movement through shared memory and registers [Thakkar_CUTLASS_2023]. The Triton backend exhibits similar trends, with rearrangement overheads further reduced and matrix multiplication dominating the total runtime.
A.4 End-to-end Prefill Speedups
@@
A.5 Post-Training Quantization Results
We compare the results of applying post-training quantization (PTQ) against Quartet using the MXFP4 format on the largest 7B model. For the PTQ baseline, we evaluate against QuaRot [ashkboos2024quarot], where the weights are quantized using GPTQ [frantar2022gptq]. To ensure a fair comparison, we introduce two key modifications to the original QuaRot approach:

### v4

@@ A.3 Scaling Law fitting
Identical to prior work [8], we fit the unquantized scaling law of the form
@@
We additionally for the scaling law forms with fixed $\gamma=1$ [27] and $\beta=1$ [28]. The fits are presented in Figure 6 alongside the mainly used of Busbridge et al. [8].
@@ A.4 Performance breakdown
Figure 7 presents a breakdown of runtime composition across three linear layer shapes in a Llama-7B model, taking the MXFP4 forward pass as an example. Each subplot shows the percentage of total runtime spent in three key kernel stages: matrix multiplication, quantization-related operations, and rearrangement of scaling factors for the mma instruction [35].
@@
As the figure illustrates, tuning the quantization kernel significantly reduces the proportion of time spent in the quantization stage—particularly for large matrix shapes. Increasing the threadblock tile size leads to more active warps per block, enhancing arithmetic intensity and enabling better latency hiding. In CUTLASS-based implementations, this change influences the multilevel tiling strategy (threadblock, warp, and instruction-level tiling), which is designed to optimize data movement through shared memory and registers [45]. The Triton backend exhibits similar trends, with rearrangement overheads further reduced and matrix multiplication dominating the total runtime.
A.5 End-to-end Prefill Speedups
@@
A.6 Post-Training Quantization Results
We compare the results of applying post-training quantization (PTQ) against Quartet using the MXFP4 format on the largest 7B model. For the PTQ baseline, we evaluate against QuaRot [2], where the weights are quantized using GPTQ [24]. To ensure a fair comparison, we introduce two key modifications to the original QuaRot approach: